I need help in writing a second equation so that the system has (-3,-3) as its only solution. The first equation in the system is y-4=3x+2. I have no idea how to solve this problem, or where to start.
Answer by Alan3354(69443)

I need help in writing a second equation so that the system has (-3,-3) as its only solution. The first equation in the system is y-4=3x+2. I have no idea how to solve this problem, or where to start.
===============
It'll be y = px + q
-3 = -3p + q
Pick a value for p, find q
eg, p = 4, q = 9
---
--> y = 4x + 9